What is the surface area of a rectangular prism with length 20 m, width 7 m, and height 3 m?
Answer by Theo(13342)

A rectangular prism has a rectangle as the base.
The area of the base is equal to length * width = 20 * 7 = 140 square cm.
The height of the rectangular prism is equal to 3 m.
The rectangular prism has 4 lateral faces.
The 3m height of the prism is equal to a lateral edge.
The area of each lateral face would be as follows:
The 2 lateral faces that connect between the 2 lengths of each base have an area of 3 * 20 = 60 square cm each.
The 2 lateral faces that connect between the 2 widths of each base have an area of 3 * 7 = 21 square cm each.
the total surface area of the rectangular prism is equal to:
2 * area of the base + 2 * area of the length connecting face + 2 * area of the width connecting face.
This is equal to 2 * 140 + 2 * 60 + 2 * 21 = 280 + 120 + 42 = 442 square cm.
***** picture not found *****